Through how many radians does the minute hand of a clock turn in 24 minutes?
Found 2 solutions by Alan3354, harpazo:
Answer by Alan3354(69443)

24/60 of 1 rev --> (24/60)*2pi radians

Answer by harpazo(655)


2*pi radians = 360 degrees
24 minutes = (24/60)*2*pi radians
(24/60) = 2/5
(2/5) * 2 * pi radians
(4pi/5) radians